OpenStudy (anonymous):

solve the given initial-value problem for y=f(x). dy/dx= (6/x)-(4/x^2) where y=2 when x=1

OpenStudy (anonymous):

separable DE: dy=(6/x -4x^-2) dx integrate... then fill in the intial conditions...

OpenStudy (anonymous):

\[\large \int\limits dy=\int\limits \frac {6}{x}-4x^{-2}dx\]

OpenStudy (anonymous):

y = 6lnx + 4/x + C Find C by plugging in y=2 when x=1

OpenStudy (anonymous):

asking for desired function i think and not a number. when i solved for c said it was incorrect